# The three phases of self-gravitating scalar field ground states

### v2

@@ Acknowledgements
We thank Noah Glennon for discussions in the beginning stages of this project. LZ would like to thank Nikhil Padmanabhan for help setting up the code that would eventually evolve to become $\mathtt{nSPIRal}$. We thank the administrative and facilities staff at the University of New Hampshire including Katie Makem-Boucher and Michelle Mancini.
@@
Appendix A Appendix
Appendix B Cosmology & Code Units
When performing numerical simulations, we use units defined according to a reference particle mass $m_{0}$:
$\mathcal{L}$ $={\left(\frac{8\pi\hbar^{2}}{3m_{0}^{2}H_{0}^{2}\Omega_{m0}}\right)}^{\frac{1}{4}}\approx 121{\left(\frac{10^{-23}\mathrm{eV}}{m_{0}}\right)}^{\frac{1}{2}}\;\mathrm{kpc},$ (22)
$\mathcal{T}$ $={\left(\frac{8\pi}{3H_{0}^{2}\Omega_{m0}}\right)}^{\frac{1}{2}}\approx 75.5\;\mathrm{Gyr},$ (23)
$\mathcal{M}$ $=\frac{1}{G}{\left(\frac{8\pi}{3H_{0}^{2}\Omega_{m0}}\right)}^{-\frac{1}{4}}{\left(\frac{\hbar}{m_{0}}\right)}^{\frac{3}{2}}\approx 7\times 10^{7}{\left(\frac{10^{-23}\mathrm{eV}}{m_{0}}\right)}^{\frac{3}{2}}M_{\odot}$ (24)
Interaction strengths are measured in units of
$\Lambda_{jk}=\frac{\hbar^{2}}{2m_{0}^{3}G\mathcal{T}}\lambda_{jk}\,.$ (25)
@@ Appendix B Cosmology & Code Units
Appendix C Imaginary Time Solver
We find spherically symmetric ground states by solving the spherically symmetric versions of the GPP equations of motion in imaginary time and using the code units described above. The steps in our imaginary time evolution obey a “kick-drift-kick” algorithm (as in Edwards et al. (2018); Glennon and Prescod-Weinstein (2021); Padmanabhan et al. (2020); Musoke (2024)). Schematically, this can be represented for $n$ iterations of timestep $h$ for wavefunction $i$ as:
$\psi_{i}(t+nh)=$ $\exp\left[-\frac{h}{2}V_{\rm{eff}}\right]$ (26)
$\times$ $\left(\prod^{n}\exp[hV_{\rm{eff}}]\exp\left[-\frac{h}{2}\nabla^{2}\right]\right)$ (27)
$\times$ $\exp\left[\frac{h}{2}V_{\rm{eff}}\right]\psi_{i}(t),\,\rm{where}$ (28)
$V_{\rm{eff}}$ $=m_{i}\Phi+\frac{1}{4}\sum_{j}m_{j}\lambda_{ij}\left|\psi_{j}\right|^{2}\psi_{i}\,.$ (29)
@@ Appendix C Imaginary Time Solver
Additionally, nSPIRal introduces a hard boundary at a given maximum radial distance $r=r_{\text{max}}$. The boundary value of each field is fixed such that $\left.r\psi_{j}\right|_{r=0}=\left.r\psi_{j}\right|_{r_{\text{max}}}=0$; the latter is equivalent to specifying $\psi_{j}(r_{\text{max}})=0$. We therefore take care to use only large enough $r_{\rm{max}}$ such that $\psi_{j}(r_{\rm{max}})\rightarrow 0$ for each field. In addition to $r_{\rm{max}}$, we must specify our grid size $n_{g}$, number of timesteps $n$, and initial conditions for the $\psi_{j}$ profiles.
Though not central to our results in this paper, we have also produced ground state profiles assuming $\lambda=0$ and plotted it in Fig. 4 in order to verify that the behavior matches predictions in the literature.

### v3

@@ Acknowledgements
We thank Noah Glennon for discussions in the beginning stages of this project and Hoang Nhan Luu for useful discussions of the manuscript. LZ would like to thank Nikhil Padmanabhan for help setting up the code that would eventually evolve to become $\mathtt{nSPIRal}$. We thank the administrative and facilities staff at the University of New Hampshire including Katie Makem-Boucher and Michelle Mancini.
@@
Appendix A Cosmology & Code Units
When performing numerical simulations, we use units defined according to a reference particle mass $m_{0}=10^{-22}$ eV. This sets the code units as follows:
$\mathcal{L}$ $={\left(\frac{8\pi\hbar^{2}}{3m_{0}^{2}H_{0}^{2}\Omega_{m0}}\right)}^{\frac{1}{4}}\approx 121{\left(\frac{10^{-23}\mathrm{eV}}{m_{0}}\right)}^{\frac{1}{2}}\;\mathrm{kpc},$ (A.1)
$\mathcal{T}$ $={\left(\frac{8\pi}{3H_{0}^{2}\Omega_{m0}}\right)}^{\frac{1}{2}}\approx 75.5\;\mathrm{Gyr},$ (A.2)
$\mathcal{M}$ $=\frac{1}{G}{\left(\frac{8\pi}{3H_{0}^{2}\Omega_{m0}}\right)}^{-\frac{1}{4}}{\left(\frac{\hbar}{m_{0}}\right)}^{\frac{3}{2}}\approx 7\times 10^{7}{\left(\frac{10^{-23}\mathrm{eV}}{m_{0}}\right)}^{\frac{3}{2}}M_{\odot}$ (A.3)
Note that the particle mass $m_{0}$ is different from the mass in the fields, $\mathcal{M}$. Interaction strengths are measured in units of
$\Lambda_{jk}=\frac{\hbar^{2}}{2m_{0}^{3}G\mathcal{T}}\lambda_{jk}\,.$ (A.4)
@@ Appendix A Cosmology & Code Units
Appendix B Imaginary Time Solver
We find nested ground states by solving the spherically symmetric versions of the GPP equations of motion in imaginary time and using the code units described above. The steps in our imaginary time evolution obey a “kick-drift-kick” algorithm (as in [63, 65, 74, 69]). Schematically, this can be represented for $n$ iterations of timestep $h$ for wavefunction $i$ as:
$\psi_{i}(t+nh)=\exp\left[-\frac{h}{2}V_{\rm{eff}}\right]\times\left(\prod^{n}\exp[hV_{\rm{eff}}]\exp\left[-\frac{h}{2}\nabla^{2}\right]\right)\times\exp\left[\frac{h}{2}V_{\rm{eff}}\right]\psi_{i}(t)\,$, (B.1)
where
$V_{\rm{eff}}=m_{i}\Phi+\frac{1}{4}\sum_{j}m_{j}\lambda_{ij}\left|\psi_{j}\right|^{2}\psi_{i}\,.$ (B.2)
@@ Appendix B Imaginary Time Solver
Additionally, nSPIRal introduces a hard boundary at a given maximum radial distance $r=r_{\text{max}}$. The boundary value of each field is fixed such that $\left.r\psi_{j}\right|_{r=0}=\left.r\psi_{j}\right|_{r_{\text{max}}}=0$; the latter is equivalent to specifying $\psi_{j}(r_{\text{max}})=0$. We therefore take care to use only large enough $r_{\rm{max}}$ such that $\psi_{j}(r_{\rm{max}})\rightarrow 0$ for each field.
In addition to $r_{\rm{max}}$, we must specify our grid size $n_{g}$, number of timesteps $n$, and initial conditions for the $\psi_{j}$ profiles. We have found that a large swatch of our parameter space is relatively insensitive to changes in these three parameters—except in vicinity of the degenerate case $m_{1}=m_{2}$, $M_{1}=M_{2}$, and $\psi_{1}(r)=\psi_{2}(r)$, as discussed in Sec. 3. Though not central to our results in this paper, we have also produced ground state profiles assuming $\lambda=0$ and plotted it in Fig. 5 in order to verify that the behavior matches predictions in the literature.